OpenStudy (anonymous):

Can Someone Explain How To Do This Please. I Will Give A Medal. :D (Solving System Equations) 2y = x + 1 –2x – y = 7

OpenStudy (anonymous):

What have you tried?

OpenStudy (anonymous):

I Havent Tried Anything Really Because I Never Understand How To Do These Problems.!

OpenStudy (abb0t):

Try substitution. Only cuz it's my favorite and I like to avoid doing a bunch of algebra :)

OpenStudy (anonymous):

Ok But What Do I Substitute.!? Im Really Confused :(

OpenStudy (abb0t):

For instance: \[ax + by = c\] and \[2dx + ey = f\] rearrange to get one variable to use to substitute for the second equation! i'll solve for y in the first one, to get: \[y = \frac{ c }{ b } - \frac{ ax }{ b }\] Now I can use that to substitue into the second equation (substitute y, because notice i have y = f(x)

OpenStudy (abb0t):

substitute all of y ino the second equation! \[2dx + e(\frac{ c }{ b }-\frac{ax}{b}) = f\]

OpenStudy (abb0t):

Notice how you have only x's on this equation now! You can combine everything into one single equation, and use algebra to solve for x. once you get a numerical value for x, plug it back into equation 1 and solve for y.

OpenStudy (anonymous):

So The I Would Put –2x – y = 7 Into The First Problem To Get y?

OpenStudy (abb0t):

Well, you have to rearrange it so you have it in the form y = f(x) You should ONLY have x's in your substitution equation.

OpenStudy (abb0t):

\[y = \frac{x}{2} + \frac{1}{2}\]

OpenStudy (anonymous):

Do I Have To Find An x Value Too? Or Whnever I Solve The Problem The x Will Be What Im Finding? So Confusing.!

OpenStudy (abb0t):

Yes. Solve for both x and y. Of course.